OpenStudy (anonymous):

put the equation into standard form. find the center the lines which contain the major and minor axes, vertices the endpoints of the minor axis, foci, and eccentricity

OpenStudy (anonymous):

\[9x^2+25y^2-54x-50y-119=0\] so i put it in order \[9x^2-54x+25y^2-50y=119\] now having a hard time completing the square

OpenStudy (anonymous):

@e.mccormick and @hartnn

OpenStudy (anonymous):

if i factor out a 9 i get and 25 i get \[9(x^2-6x+9)+25(y^2-10y+25)=119\] is this right so far ?

OpenStudy (e.mccormick):

Hmmm.... how did you get the y part?

OpenStudy (anonymous):

oops should be a 2 not 25

OpenStudy (anonymous):

damn here \[9(x^2-6x+9)+25(y^2-2y+1)=119\] ?

OpenStudy (anonymous):

so end will end up being 225? then divide by 225 guess i just needed to talk it through

OpenStudy (e.mccormick):

You also have to ballance what you add with something you subtract. I leave things in ( ) to make it a little clearer what I am doing with the sign, but it is optional. \(9(x^2-6x+(-3)^2 - (-3)^2)+25(y^2-2y +(-1)^2-(-1)^2)=119\)

OpenStudy (anonymous):

thanks i got it :) just guess i needed to talk it out haha

OpenStudy (e.mccormick):

Yah, and do you know what that simplifies to?

OpenStudy (anonymous):

its \[\frac{ (x-3)^2 }{ 25 }+\frac{ (y-1)^2 }{ 9 }=1\]

OpenStudy (e.mccormick):

The \(+(-3)^2\) becomes part of the square. The \(-(-3)^2\) gets evaluated. Same sort of thing on the y part. \(9((x-3)^2 - 9)+25((y-1)^2-1)=119\) I did not do the rest yet.

OpenStudy (e.mccormick):

\((x-3)^2 - 81 +(y-1)^2-25=119\) \((x-3)^2 +(y-1)^2=119+ 81+ 25\) \((x-3)^2 +(y-1)^2=225\) Yah, looks like it.

OpenStudy (e.mccormick):

Oops. missed typing in the 9 and 25. LOL.

OpenStudy (anonymous):

:)

OpenStudy (e.mccormick):

I meant: \(9(x-3)^2 - 81 +25(y-1)^2-25=119\) \(9(x-3)^2 +25(y-1)^2=119+ 81+ 25\) \(9(x-3)^2 +25(y-1)^2=225\) =P

OpenStudy (e.mccormick):

Yah, so: \(\dfrac{ (x-3)^2 }{ 5^2 }+\dfrac{ (y-1)^2 }{ 3^2 }=1\)